Two conductors of se radii r and 2r are connected in 1) series 2) in parallel. Find the ratio of their drift velocities in each case.

Grade:12

1 Answers

Khimraj
5 years ago
​i) as current will be same
 so I = n*e*A1*vd1 =n*e*A2​*vd2
so vd1/vd2 = r22/r12 = 4/1
Hope it clears.